OpenStudy (anonymous):

∫(0,2)∫(y,4) sin(x)√xdxdy

OpenStudy (anonymous):

what are you trying to do here?

OpenStudy (anonymous):

evaluate the double integral

OpenStudy (anonymous):

But I have to switch the order and i dont know how to do that.

OpenStudy (anonymous):

is this like this?? \[\int\limits_{0}^{2}\int\limits_{y}^{4}\sin(x)\sqrt{x}dxdy\]

OpenStudy (anonymous):

yes but the integral for dx goes from y^2 to 4 sorry

OpenStudy (anonymous):

its equal to sin4 - 4cos4 i can show steps but it won't be till later today.

OpenStudy (anonymous):

could you just tell me the integral then after the order has been switched!!

OpenStudy (ash2326):

@EEmarr Sorry but I have trouble recalling the methodology used here, do you have any idea how to proceed? maybe then I can contribute something

OpenStudy (anonymous):

I don't think EEmarr is online. But my teacher said this integral would be impossible if i dont switch the order of integration.

OpenStudy (anonymous):

switching the order has to do with looking at where the graph lays and the limits that are already being used. you then have to interchange the x's and y's. i haven't worked these problems in over a year so it will take me a while to look over the material.

OpenStudy (anonymous):

ok so the limits should change to \[\int\limits_{0}^{4}\int\limits_{0}^{\sqrt{x}}\] look at this video is explains it faster than i can. also on the first one it may be from 0 to 2 i can't tell without drawing a very detailed graph.

OpenStudy (anonymous):

http://www.youtube.com/watch?v=NETmfwOAKpQ

OpenStudy (anonymous):

by first one, do you mean the integral for dx or dy?

OpenStudy (anonymous):

I will check out the video, thanks!

OpenStudy (anonymous):

so when you change the order from dxdy. it will then be dydx so i mean dy

OpenStudy (anonymous):

Ok, by the way would i still be integrating sqrt(x)sinx

OpenStudy (anonymous):

All right, I got the answer. Thank You very much!!!!!

OpenStudy (anonymous):

No problem